A biker bikes 1450 meters uphill in 15 minutes. He averages 2.4 m/s for the entire 2720 meter portion of the trip. What is his average velocity on the way down the hill.
1
Expert's answer
2020-09-28T08:08:39-0400

The average velocity is total distance divided by elapsed time:


"v=2.4=\frac{D}{T}=\frac{2720}{15\text{ min}\cdot60\text{ s/min}+t_\text{down}},\\\space\\
t_\text{down}=233\text{ s}."

The average velocity down is


"v_\text{down}=\frac{d_\text{down}}{t_\text{down}}=\frac{2720-1450}{233}=5.45\text{ m/s}."